OpenStudy (hijacktrolo):

Find an equivalent function to f(x) = 3(6)2x. f(x) = 3(36)x f(x) = 182x f(x) = 108x f(x) = 9x(36x) I think the answer is c

OpenStudy (hijacktrolo):

hang on let me clarify some things

OpenStudy (hijacktrolo):

ok so its to the power of 2x

OpenStudy (hijacktrolo):

f(x) = 3(36)^x f(x) = 18^2x f(x) = 108^x f(x) = 9^x(36^x)

OpenStudy (hijacktrolo):

f(x) = 3(6)^2x

OpenStudy (s4sensitiveandshy):

\[f(x)=3(6)^{2x}\] f(x)=3(6)^(2x) parenthess ^

OpenStudy (hijacktrolo):

the first way you said it was right

OpenStudy (s4sensitiveandshy):

\[f(x)=3(6)^{2x}\] can be written as \[f(x)=3(6^2)^x\] exponent rule: \[(a)^{nx} = (a^n)^x\]

OpenStudy (hijacktrolo):

so c :/

OpenStudy (s4sensitiveandshy):

what's C ? is it f(x) = (108)^x ??

OpenStudy (hijacktrolo):

yes

OpenStudy (s4sensitiveandshy):

no that's not correct. you can't multiply 3 with (6^2)^x because (6^2) is raised to the x \[4 * 9^x \cancel{=} 36^x\]

OpenStudy (hijacktrolo):

so the answer would be d?

OpenStudy (hijacktrolo):

im confused lol

OpenStudy (s4sensitiveandshy):

\[f(x)=3(6)^{2x}\] like i said we can rewrite \[6^{2x} as (6^2)^x\] \[f(x)=3(6^2)^x\] now simplify the equation

OpenStudy (hijacktrolo):

f(x)=3(36)^x?

OpenStudy (mathmale):

Please make a habit of using the character " ^ " to denote exponentiation. We've already lost time here by having to verify your meaning. f(x) = 3(6)2x would be best expressed by \[f(x)=3(6)^{2x}\] if you're willing to use Equation Editor. Otherwise use parentheses around the exponent: 3(6)^(2x). sensitiveandshy is right on target in her explanations. Yes, your answer is correct. Nice work!

OpenStudy (s4sensitiveandshy):

let me retype all the options \[f(x)=3(36)^x\]\[f(x)=(18)^{2x}\]( i think 2nd option looks like this ) \[f(x)=(108)^x\] Don't know if it is 9 times (36x) or 9^x(36)^x

OpenStudy (s4sensitiveandshy):

yes that's correct.

OpenStudy (hijacktrolo):

:D thank both and ill make sure to do that from now on

OpenStudy (mathmale):

Good going! Thank you for being open to suggestions.